The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught. Step 1 of 2 : Suppose a sample of 716 suspected criminals is drawn. Of these people, 286 were captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution :

Given that n = 716 , x = 286

=> proportion p = x/n

= 286/716

= 0.3994

= 0.399 (rounded)

=> The proportion of people who were caught after being on the 10 Most Wanted list is 0.399
Answer #2

To estimate the proportion of people who were caught after being on the 10 Most Wanted list, we can use the following formula:

Proportion = (Number of captured individuals) / (Total sample size)

Given that a sample of 716 suspected criminals was drawn, and 286 of them were captured, we can calculate the proportion as follows:

Proportion = 286 / 716

Proportion ≈ 0.399

Rounded to three decimal places, the estimated proportion of people who were caught after being on the 10 Most Wanted list is approximately 0.399.

answered by: Hydra Master